OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. (8,-7) ; 4x-7y=8

OpenStudy (anonymous):

thank you for helping me

OpenStudy (amistre64):

since the slope there is -4/-7; we flip that number and negate it -7/4 works great 7x+4y = b 7(8) +4(-7) = b 56 -28 = b 28 = b 7x +4y = 28 is the equation perpendicular.... you wanted parallel tho ....back to the drawing board..

OpenStudy (amistre64):

4(8) -7(-7) = b 32 +49 = b 81 = b 4x -7y = 81 is the parallel line :)

OpenStudy (anonymous):

Thanks for your help :)

OpenStudy (amistre64):

youre welcome :)